OpenStudy (anonymous):

Write the quadratic function with the following transformations: vertical shrink by 3, left 4, down 9

OpenStudy (anonymous):

\[f(x)=\frac{ 1 }{ 3 }(x-4)^{2}-9\]

OpenStudy (anonymous):

should be x+4

OpenStudy (loser66):

yes,

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

\[f(x)=\frac{ 1 }{ 3 }(x+4)^{2}-9\]

OpenStudy (mathstudent55):

f(x) = a(x - h)^2 + k a = width = vertical stretch h = horizontal shift k = vertical shift

OpenStudy (mathstudent55):

a between 0 and 1 is vertical shrink, so 1/3 is correct.

OpenStudy (mathstudent55):

vertical shrink means parabola becomes wider.

OpenStudy (jhannybean):

Ah ok.i got it.